# Metric Rarity and the Rise of Symmetry in Invariant Optimization

### v1

@@ 1.2.2 Regime I: The Empty Interior (Ubiquitous Symmetry)
The geometric rarity established above is not a static global constraint; rather, it operates recursively across the lattice of subgroups. Consider the search for critical points possessing a specific symmetry group $H\subseteq G$. By the Principle of Symmetric Criticality (Theorem 2.6), this problem reduces to finding critical points of the restriction $f|_{X(\mathbb{R})^{H}}$ on the fixed-point subspace.
@@
This relation uncovers a cascading selection mechanism. For small subgroups $H$, the effective group $\overline{G}$ remains large (comparable to $S_{n}$), resulting in a massive number of involutions and a metrically rare real image $L_{H}$. In this regime, finding a critical point is statistically improbable. However, as we restrict our attention to strata with larger stabilizers $H$, the effective group $\overline{G}$ shrinks, the number of involutions $\mathrm{Inv}(\overline{G})$ decreases, and the relative volume of the real image increases. This monotonicity implies that the system is statistically driven away from low-symmetry strata and ”settles” only when the symmetry $H$ is sufficiently high to render the real image $L_{H}$ voluminous enough to capture the critical points of the potential.
In Section 6, we formalize the connection between this theoretical mechanism and the empirical data. We translate the results of [1, 2, 3] into this algebraic language, demonstrating how the ”Empty Interior” phenomenon provides an explanation for the experimentally observed ubiquity of symmetry.
@@ 1.2.3 Regime II: The Active Constraint (A Geometric Hypothesis)
While the metric rarity established in Regime I successfully explains the baseline ubiquity of symmetry, it leaves a fundamental phenomenon unaccounted for: the energetic ordering. The metric argument implies that critical points should settle into the first stratum $L_{H}$ that is voluminous enough within $Y^{(H)}(\mathbb{R})$ to statistically accommodate them. However, this fails to explain why, across both random $S_{n}$-invariant polynomials and physical clusters (e.g., $LJ_{13}$), the deepest minima exhibit significantly higher symmetry than the bulk critical points.
This is particularly striking in $LJ_{13}$, where the “Empty Interior” condition does not hold -approximately 94% of critical points and approximately 78% of the local minima are asymmetric—yet the system bypasses these traps to locate the unique, maximally symmetric icosahedron at the global minimum.
To address this, we propose a second geometric mechanism: the Active Constraint.
The Carpet and the Himalayas.
We envision the landscape $\tilde{f}\mid_{L\subset Y(\mathbb{R})}$ through a geometric analogy: consider a carpet $L$ cast arbitrarily upon the vast, rugged terrain of the Himalayas (representing the ambient potential $\tilde{f}$ on $Y(\mathbb{R})$). In this picture, while the carpet may have many small local wrinkles (local minima), it is dominated by the massive global slope of the underlying mountains.
Just as the lowest point of such a carpet is almost certain to lie on its edge—and specifically at a sharp corner—rather than in a central hollow, we propose that the global minimum of $\tilde{f}|_{L}$ is driven to the boundary $\partial L$. In our algebraic dictionary, the boundary $\partial L$ is the locus of non-trivial stabilizers (see Remark 3.3), and ”sharper corners” correspond to higher symmetry strata. Thus, the global minimum is attained at a corner because it is there that the global descent is finally intercepted by the domain’s geometry.
@@
From our perspective, a ”funnel” is the physical manifestation of a coherent global gradient of $\tilde{f}$ across the domain $L\subset Y(\mathbb{R})$. Even when the landscape $\tilde{f}$ is rugged and filled with local minima, the essential geometric feature remains: these gradients do not vanish in the interior. Instead, they create a persistent downhill drift that directs the system ”outwards” effectively pinning the global minimum against the ”walls” $\partial L$. The regular icosahedron of $LJ_{13}$ is thus the ”sharp corner” of the real image LL—the point where the global descent of the potential $\tilde{f}\mid_{Y(\mathbb{R})}$ reaches the limit of the domain $L$.
@@ 1.3 Related Work
The question of the origin of symmetry in physical systems has engaged researchers since the earliest days of mathematical physics. As early as 1894, Pierre Curie proposed the guiding principle that “the symmetry of the effect is the maximal symmetry compatible with the existence of the phenomenon” [6]. In the context of discrete geometry and optimization, Cohn [7] reframes this challenge as the search for the “genetics of regular figures”—a term coined by Fejes Tóth to describe the economy principle that generates order out of chaotic sets. Cohn argues that while our algebraic classification of symmetry groups is advanced, our understanding of this generative mechanism still “lags behind.”
@@
This search for a “genetic” mechanism is motivated by robust empirical evidence indicating that complex systems actively prefer symmetry. In chemical physics, the “Epikernel Principle” formulated by Ceulemans and Vanquickenborne [11] suggests that when a symmetric system becomes unstable (e.g., via the Jahn-Teller effect), symmetry breaking is not total. Instead, the system prefers to descend into maximal subgroups (epikernels), effectively “sticking” to high-symmetry strata rather than collapsing to complete asymmetry.
@@
While explicit counterexamples have demonstrated that Michel’s conjecture is not universally true [18], the strong statistical tendency for high-symmetry minima remains a defining feature of these landscapes. Crucially, a distinction in scope exists between these approaches and the present work. While the Michel-Abud-Sartori program focuses on defining the strict conditions under which maximal symmetry is deterministically guaranteed, our framework utilizes the metric rarity of the real image to explain the statistical preference for high symmetry (not necessarily maximal). We provide a mechanism that operates generically across broad families of groups and potentials, outside the specific constraints required for the validity of Michel’s Conjecture.
This preference for order extends to the thermodynamic limit ($N\to\infty$). As reviewed by Blanc and Lewin [19], the “Crystallization Conjecture” posits that interacting particles spontaneously arrange in periodic structures at low temperatures. While rigorous proofs—such as Theil’s for Lennard-Jones-like potentials [20]—are rare and analytically demanding, the empirical ubiquity of crystallization underscores the universality of symmetry selection.
@@
Despite these significant theoretical and empirical advances, a comprehensive answer to the fundamental question of symmetry selection remains elusive. While the existing frameworks successfully address specific aspects—ensuring the existence of symmetric critical points, verifying their local stability, or rationalizing their statistical presence in the tails of distributions—none offers a complete, unified explanation for the pervasive dominance of symmetry in global optimization. The mechanism by which a complex system effectively “maps” the vast asymmetric landscape to locate a rare symmetric minimum remains, to a large extent, obscure. In this work, we aim to bridge this gap by constructing a rigorous deterministic mechanism that serves as the geometric substrate for this statistical tendency.
@@ 1.4 Paper Organization
In Section 2, we establish the algebraic foundation, defining the general $G$-variety setup, the quotient space construction, and the Principle of Symmetric Criticality. Section 3 provides the geometric dictionary connecting algebra to topology. We prove that the differential of the quotient map is surjective if and only if the stabilizer is trivial (Theorem 3.1) and identify configurations with even order stabilizer with the boundary of the real image (Theorem 3.2). In Section 4, we quantify the ”metric rarity” of the real image. We begin by analyzing the quotient of $S_{n}$ acting on $\mathbb{R}^{n}$, utilizing probabilistic results regarding polynomials with real roots. We then define a natural measure for general finite group actions on vector spaces to calculate the rarity of the real image, and discuss how this rarity applies recursively across fixed-point strata. Section 5 extends this framework to physical particle systems. We formalize the Shape Space and the Total Quotient, proving that the volume of physically realizable configurations decays exponentially within the ambient quotient (Theorem 5.13). In Section 6, we apply our theoretical framework to interpret empirical observations. We map the experimental setups from the works of Arjevani et al. and Schneider to our algebraic model, demonstrating how the metric rarity of the real image explains the ubiquity of symmetry (”Regime I”) in these landscapes. Finally, Section 7 summarizes the findings and discusses the implications for Energy Landscape Theory and future research directions.
@@ 2 The Algebraic Framework
This principle is fundamental to our analysis: it establishes that symmetric critical points are stable features of the landscape rather than artifacts. However, while Theorem 2.6 guarantees their existence, it implies nothing about their abundance.
@@ 3 The Geometry of the Quotient Map: Singularities and Boundaries
To understand why physical potentials ”prefer” symmetric configurations, we analyze the geometry of the real image $L=\pi(X(\mathbb{R}))$. In this section, we establish a fundamental dictionary between the algebraic properties of a point (its stabilizer) and its topological placement within the quotient (interior vs. boundary).
@@
Theorem 3.1 (Surjectivity of the Quotient Differential).
@@
Since $G$ is finite and $x\in X(\mathbb{R})$ is a smooth point, we apply Luna’s Slice Theorem [24, Theorem 5.3]. Let $V=T_{x}X$ be the tangent space at $x$. Let $V(\mathbb{R})$ denote the set of real points of this vector space.
According to the Slice Theorem, there exists a $G_{x}$-invariant locally closed subvariety $S\subset X$ containing $x$, such that the natural $G$-morphism $\psi:G\times^{G_{x}}S\to X$ is excellent (see [24, Definition 5.7]). Since $x$ is a smooth point, we can identify $S$ with a $G_{x}$-invariant neighborhood $U$ of 0 in $V$.
@@
Here, $q$ is induced by the projection to the quotient of the linear action. Since $\overline{\psi}$ is an étale morphism of real varieties (a property implied by $\psi$ being excellent [24, Definition 5.7(i)]), it induces a local homeomorphism on the set of real points.
The horizontal maps $\psi$ and $\overline{\psi}$ are étale [24, Theorem 5.3]. Étale morphisms induce isomorphisms on the tangent spaces. Therefore, the surjectivity of $d\pi_{x}$ is equivalent to the surjectivity of the differential of the quotient map $dq_{0}:T_{0}V\to T_{q(0)}(V/\!/G_{x})$ at the origin.
@@
Theorem 3.2 (The Boundary of the Real Image and Stabilizer Parity).
@@ 4.1 Metric Rarity in the Symmetric Group Quotient
To quantify the ”rarity” of the real image, we first focus on the canonical case of the symmetric group $S_{n}$ acting on the configuration space $\mathbb{R}^{n}$. In this setting, the quotient map $\pi:\mathbb{C}^{n}\to\mathbb{C}^{n}/S_{n}\cong\mathbb{C}^{n}$ is realized by the elementary symmetric polynomials (Vieta’s map):
@@
We now demonstrate that for large $n$, the ”island” of real-rooted polynomials $L$ becomes rare within the ”ocean” of generic monic polynomials $Y(\mathbb{R})$.
@@
Let $\mu_{n}=\frac{1}{n}\sum_{z:Q_{n}(z)=0}\delta_{z}$ denote the empirical measure of the roots of $Q_{n}$. We invoke the Large Deviation Principle (LDP) established by Butez [21, Theorem 1.5] (extending Zeitouni–Zelditch [22]). The sequence of random measures $(\mu_{n})$ satisfies an LDP in the space of probability measures $\mathcal{M}_{1}(\mathbb{C})$ with speed $\beta_{n}=n^{2}$ and a good rate function $\tilde{I}$.
Crucially, the rate function $\tilde{I}(\mu)$ possesses a unique minimizer: the uniform measure on the unit circle $\nu_{S^{1}}$ (see [21], Section 1). The event that all roots are real corresponds to the set of measures supported on the real line:
@@
We apply this geometric decay to the statistical analysis of critical points. Let $f$ be an $S_{n}$-invariant polynomial that descends to a polynomial $\tilde{f}:Y(\mathbb{R})\to\mathbb{R}$ via the relation $f=\tilde{f}\circ\pi$. If $\tilde{f}$ is a polynomial of degree $n$, then by Bézout’s Theorem, its gradient system $\nabla\tilde{f}=0$ possesses at most $n^{n}$ complex critical points. Recall from Theorem 3.1 that an asymmetric critical point of $f$ corresponds uniquely to a critical point of $\tilde{f}$ lying in the smooth locus of the real image $L^{\circ}$. This reduces the question to: what is the probability that any of these $n^{n}$ candidates falls into the domain $L$?

### v2

@@ 1.2.2 Regime I: The “Empty Interior” (No Interior Critical Points)
The geometric rarity established above extends recursively across the lattice of subgroups. Consider the search for critical points possessing a specific symmetry group $H\subseteq G$. By the Principle of Symmetric Criticality (Theorem 2.6), this problem reduces to finding critical points of the restriction $f|_{X(\mathbb{R})^{H}}$ on the fixed-point subspace.
@@
This relation uncovers a cascading selection mechanism. For small subgroups $H$, the effective group $\overline{G}$ remains large (comparable to $S_{n}$), resulting in a large number of involutions and a metrically rare real image $L_{H}$. In this regime, finding a critical point is statistically improbable. However, as we restrict our attention to strata with larger stabilizers $H$, the effective group $\overline{G}$ shrinks, the number of involutions $\mathrm{Inv}(\overline{G})$ decreases, and the relative volume of the real image increases. This monotonicity implies that the system is statistically driven away from low-symmetry strata and resides only where the symmetry $H$ is sufficiently high to render the real image $L_{H}$ voluminous enough to capture the critical points of the potential.
In Section 6, we formalize the connection between this theoretical mechanism and the empirical data. We translate the results of [1, 2, 3] into this algebraic language, demonstrating how the “Empty Interior” phenomenon (no critical points of $\tilde{f}$ in $L^{\circ}$) provides an explanation for the experimentally observed prevalence of symmetry.
@@ 1.2.3 Regime II: The Active Constraint (A Geometric Hypothesis)
While the metric rarity established in Regime I successfully explains the baseline prevalence of symmetry, it leaves a fundamental phenomenon unaccounted for: the energetic ordering. The metric argument implies that critical points should reside in the first stratum $L_{H}$ that is voluminous enough within $Y^{(H)}(\mathbb{R})$ to statistically accommodate them. However, this fails to explain why, across both random $S_{n}$-invariant polynomials and physical clusters (e.g., $LJ_{13}$), the deepest minima exhibit higher symmetry than the bulk critical points.
In $LJ_{13}$, the “Empty Interior” condition does not hold—approximately 94% of critical points and 78% of local minima are asymmetric. Nevertheless, the landscape is governed by a strict energetic ordering: the global minimum is attained at the unique, maximally symmetric icosahedron. Furthermore, this is not an outlier; the density of non-trivial symmetries increases monotonically as energy decreases, effectively localizing the deepest critical points within the high-symmetry strata of the boundary $\partial L$. To address this, we propose a second geometric mechanism: the Active Constraint.
Heuristic description.
We visualize the landscape of $\tilde{f}|_{L}$ through a geometric analogy: consider a bounded domain $L$ (analogous to a “small carpet”) placed upon a vast, rugged terrain (analogous to the “Himalayas”). If the domain were large, it might capture multiple deep valleys, potentially containing the global minimum in its interior. However, due to the metric rarity established above, $L$ is effectively small relative to the ambient landscape’s variation.
Consequently, the domain is likely to rest upon a single dominant slope. While local non-convexities may create shallow interior minima, the overarching ambient gradient drives the global minimum toward the boundary $\partial L$, where the descent is geometrically arrested. In our algebraic framework, the boundary $\partial L$ is the locus of non-trivial stabilizers (see Remark 3.3), and strata of higher codimension correspond to larger stabilizers. Thus, the global minimum is driven to a high-codimension stratum of $\partial L$, where the global descent is intercepted by the geometry of the domain.
@@
From our perspective, a ”funnel” can be understood as the physical manifestation of a coherent global gradient of $\tilde{f}$ across the domain $L\subset Y(\mathbb{R})$. Even when the landscape $\tilde{f}$ is multimodal, the essential geometric feature remains: the ”global gradient” does not vanish in the interior. Instead, it creates a persistent downhill drift that directs the system outward, constraining the global minimum to the boundary $\partial L$. In this framework, the regular icosahedron of $LJ_{13}$ corresponds to a high-codimension stratum of the real image LL—the locus where the global descent of the potential $\tilde{f}\mid_{Y(\mathbb{R})}$ reaches the boundary of the domain $L$.
@@
1.3 Statement of Contributions: Results and Proposed Mechanisms
We distinguish explicitly between rigorous geometric results, empirical findings, and the mechanisms proposed to explain them.
I. Mathematical Results.
We consider the general setting where $X$ is an irreducible complex affine algebraic variety defined over $\mathbb{R}$, equipped with a faithful action of a group $G$. Within this framework, we establish the following theorems:
  1. 1.
    Probabilistic Rarity (The Gaussian $S_{n}$ Case) (Theorem 4.2): Specializing to $G=S_{n}$, we consider the standard Gaussian measure directly on the quotient space $Y(\mathbb{R})\cong\mathbb{R}^{n}$. Proceeding under Assumption 4.1 (the asymptotic equivalence of Monic and Kac ensembles), we deduce that the measure of the real image $L$ decays super-exponentially as $\sim e^{-Cn^{2}}$, rendering the asymmetric stratum statistically small for large $n$.
  2. 2.
    Geometric Rarity (General Finite Groups) (Theorem 4.5): For finite groups acting on vector spaces, we introduce a natural measure on the quotient $Y(\mathbb{R})$ induced by a $G$-invariant Hermitian metric on $X$. Within this framework, we prove that the relative volume of the real image $L$ is exactly $(\#\mathrm{Inv}(G))^{-1}$, identifying the number of involutions as the controlling parameter for the scarcity of the physical domain.
  3. 3.
    Extension to Reductive Groups (Particle Systems) (Theorem 5.13): We extend the metric framework to the reductive group action governing particle systems, $G=O(d,\mathbb{C})\times S_{n}$. Using a Hermitian metric invariant under the maximal compact subgroup $K=S_{n}\times O(d,\mathbb{R})$, we prove that the physically realizable domain $L$ constitutes a vanishingly small fraction of the ambient quotient, with relative volume decaying as $\sim 2^{-\min(d,n)}\cdot(\#\mathrm{Inv}(S_{n}))^{-1}$.
Unified Algebraic Perspective. Beyond these specific estimates, we observe (Remark 5.16) that the volume constraints derived above can be unified via Galois Cohomology: the physical image corresponds to the trivial cohomology class, while the ambient space is populated by non-trivial real forms.
II. Empirical Findings.
Complementing the theoretical analysis, we report the following experimental result:
  • Universality of Energetic Ordering (Regime II): We document extensive numerical evidence (Appendix B) showing that the “Regime II” phenomenon—where symmetry increases as energy decreases—is not limited to physical potentials. We establish that this monotonic ordering emerges generically in random $S_{n}$-invariant polynomial landscapes, identifying it as a structural feature of invariant optimization rather than a consequence of specific physical forces.
III. Proposed Mechanisms.
Based on the geometric rarity and the empirical ordering, we propose two driving mechanisms:
  1. 1.
    Regime I: The “Empty Interior” (No Critical Points in $L^{\circ}$). For non-convex potentials of moderate complexity (where the number of critical points does not overwhelm the super-exponential volume decay of $\mu(L)$), we posit that critical points are statistically excluded from the smooth locus of $L^{\circ}$. This forces critical points to reside on the singular locus, resulting in prevalent symmetry.
  2. 2.
    Regime II (Energetic ordering by symmetry). We formulate the Active Constraint hypothesis: due to the metric rarity of the real image $L$, the physical domain acts as a small, bounded manifold subject to the ambient potential $\tilde{f}$. Consequently, the landscape restricted to $L$ is dominated by a coherent global gradient rather than local fluctuations. This structural configuration localizes the lowest energy states within the high-codimension strata of the boundary $\partial L$. We interpret the “funnel” topography of Lennard-Jones clusters simply as the manifestation of such a global gradient oriented toward the boundary strata of high symmetry.
@@ 1.4 Related Work
The question of the origin of symmetry in physical systems has engaged researchers since the earliest days of mathematical physics. As early as 1894, Pierre Curie proposed the guiding principle that “the symmetry of the effect is the maximal symmetry compatible with the existence of the phenomenon” [8]. In the context of discrete geometry and optimization, Cohn [9] reframes this challenge as the search for the “genetics of regular figures”—a term coined by Fejes Tóth to describe the economy principle that generates order out of chaotic sets. Cohn observes that while the algebraic classification of symmetry groups is well developed, a satisfactory explanation of this generative mechanism remains open.
@@
This search for a “genetic” mechanism is motivated by robust empirical evidence indicating that complex systems exhibit a statistical preference for symmetry. In chemical physics, the “Epikernel Principle” formulated by Ceulemans and Vanquickenborne [13] suggests that when a symmetric system becomes unstable (e.g., via the Jahn-Teller effect), symmetry breaking is not total. Instead, the system prefers to descend into maximal subgroups (epikernels), localizing in high-symmetry strata rather than collapsing to complete asymmetry.
@@
Similarly, in the context of discrete geometry—primarily concerning point configurations on the sphere—Cohn and Kumar [19] formulated the rigorous concept of universal optimality. Consider a configuration of $N$ points $x=(x_{1},\dots,x_{N})$ on the unit sphere $S^{d-1}$. The total potential energy is defined as a functional $E_{f}:(S^{d-1})^{N}\to\mathbb{R}$ given by the pairwise sum:
$E_{f}(x)=\sum_{1\leq i<j\leq N}f(\|x_{i}-x_{j}\|^{2}).$
Universal optimality requires a configuration to minimize $E_{f}$ globally, simultaneously for the entire class of completely monotonic interaction functions.
While explicit counterexamples refute the universality of Michel’s conjecture [21], and universal optima are guaranteed only in rare cases, the statistical tendency for high-symmetry minima persists. A key distinction in scope exists between these approaches and the present work. While the Michel-Abud-Sartori program and Cohn’s framework focus on deterministic guarantees for maximal or universal symmetry, our framework utilizes the metric rarity of the real image to explain the statistical preference for high (not necessarily maximal) symmetry. This mechanism operates generically across broad families of potentials, outside the strict constraints required for universal optimality.
This preference for order extends to the thermodynamic limit ($N\to\infty$). As reviewed by Blanc and Lewin [22], the “Crystallization Conjecture” posits that interacting particles spontaneously arrange in periodic structures at low temperatures. While rigorous proofs—such as Theil’s for Lennard-Jones-like potentials [23]—are rare and analytically demanding, the empirical prevalence of crystallization underscores the universality of symmetry selection.
@@
Despite these significant theoretical and empirical advances, a comprehensive answer to the fundamental question of symmetry selection remains elusive. While the existing frameworks successfully address specific aspects—ensuring the existence of symmetric critical points, verifying their local stability, or rationalizing their statistical presence in the tails of distributions—none offers a complete, unified explanation for the prevalence of symmetric critical points in global optimization. The mechanism by which a complex system selects a symmetric global minimum from a configuration space in which symmetric points have measure zero remains not fully understood. In this work, we aim to bridge this gap by constructing a rigorous deterministic mechanism that serves as the geometric substrate for this statistical tendency.
@@ 1.5 Paper Organization
In Section 2, we establish the algebraic foundation, defining the general $G$-variety setup, the quotient space construction, and the Principle of Symmetric Criticality. Section 3 provides the geometric dictionary connecting algebra to topology. We prove that the differential of the quotient map is surjective if and only if the stabilizer is trivial (Theorem 3.1) and identify configurations with even order stabilizer with the boundary of the real image (Theorem 3.2). In Section 4, we quantify the ”metric rarity” of the real image. We begin by analyzing the quotient of $S_{n}$ acting on $\mathbb{R}^{n}$, utilizing probabilistic results regarding polynomials with real roots. We then define a natural measure for general finite group actions on vector spaces to calculate the rarity of the real image, and discuss how this rarity applies recursively across fixed-point strata. Section 5 extends this framework to physical particle systems. We formalize the Shape Space and the Total Quotient, proving that the volume of physically realizable configurations decays exponentially within the ambient quotient (Theorem 5.13). In Section 6, we apply our theoretical framework to interpret empirical observations. We map the experimental setups from the works of Arjevani et al. and Schneider to our algebraic model, demonstrating how the metric rarity of the real image explains the prevalence of symmetry (“Regime I”) in these landscapes. Finally, Section 7 summarizes the findings and discusses the implications for Energy Landscape Theory and future research directions.
@@ 2 The Algebraic Framework
Theorem 2.6 guarantees the existence of symmetric critical points but implies nothing about their abundance.
@@ 3 The Geometry of the Quotient Map: Singularities and Boundaries
The statistical preference of invariant potentials for symmetric configurations can be traced to the geometry of the real image $L=\pi(X(\mathbb{R}))$. In this section, we establish a correspondence between the algebraic properties of a point (its stabilizer) and its topological placement within the quotient (interior vs. boundary).
@@
Proposition 3.1 (Surjectivity of the Quotient Differential; see e.g., [28]).
@@
Since $G$ is finite and $x\in X(\mathbb{R})$ is a smooth point, we apply Luna’s Slice Theorem [28, Theorem 5.3]. Let $V=T_{x}X$ be the tangent space at $x$. Let $V(\mathbb{R})$ denote the set of real points of this vector space.
According to the Slice Theorem, there exists a $G_{x}$-invariant locally closed subvariety $S\subset X$ containing $x$, such that the natural $G$-morphism $\psi:G\times^{G_{x}}S\to X$ is excellent (see [28, Definition 5.7]). Since $x$ is a smooth point, we can identify $S$ with a $G_{x}$-invariant neighborhood $U$ of 0 in $V$.
@@
Here, $q$ is induced by the projection to the quotient of the linear action. Since $\overline{\psi}$ is an étale morphism of real varieties (a property implied by $\psi$ being excellent [28, Definition 5.7(i)]), it induces a local homeomorphism on the set of real points.
The horizontal maps $\psi$ and $\overline{\psi}$ are étale [28, Theorem 5.3]. Étale morphisms induce isomorphisms on the tangent spaces. Therefore, the surjectivity of $d\pi_{x}$ is equivalent to the surjectivity of the differential of the quotient map $dq_{0}:T_{0}V\to T_{q(0)}(V/\!/G_{x})$ at the origin.
@@
Theorem 3.2 (Boundary Stratification; adapted from [7]).
@@ 4.1 Metric Rarity in the Symmetric Group Quotient
We begin by quantifying the measure of the real image in the canonical case of the symmetric group $S_{n}$ acting on the configuration space $\mathbb{R}^{n}$. In this setting, the quotient map $\pi:\mathbb{C}^{n}\to\mathbb{C}^{n}/S_{n}\cong\mathbb{C}^{n}$ is realized by the elementary symmetric polynomials (Vieta’s map):
@@
We now demonstrate that for large $n$, the subset $L$ of real-rooted polynomials has vanishing measure relative to the full coefficient space $Y(\mathbb{R})$.
@@
Let $\mu_{n}=\frac{1}{n}\sum_{z:Q_{n}(z)=0}\delta_{z}$ denote the empirical measure of the roots of $Q_{n}$. We invoke the Large Deviation Principle (LDP) established by Butez [24, Theorem 1.5] (extending Zeitouni–Zelditch [25]). The sequence of random measures $(\mu_{n})$ satisfies an LDP in the space of probability measures $\mathcal{M}_{1}(\mathbb{C})$ with speed $\beta_{n}=n^{2}$ and a good rate function $\tilde{I}$.
The rate function $\tilde{I}(\mu)$ possesses a unique minimizer: the uniform measure on the unit circle $\nu_{S^{1}}$ (see [24], Section 1). The event that all roots are real corresponds to the set of measures supported on the real line:
@@
We apply this geometric decay to the statistical analysis of critical points. Let $f$ be an $S_{n}$-invariant polynomial that descends to a polynomial $\tilde{f}:Y(\mathbb{R})\to\mathbb{R}$ via the relation $f=\tilde{f}\circ\pi$. If $\tilde{f}$ is a polynomial of degree $n$, then by Bézout’s Theorem, its gradient system $\nabla\tilde{f}=0$ possesses at most $n^{n}$ complex critical points. Recall from Theorem 3.1 that an asymmetric critical point of $f$ corresponds uniquely to a critical point of $\tilde{f}$ lying in the smooth locus of the real image $L^{\circ}$. The problem thus reduces to estimating the probability that any of these $n^{n}$ candidates lies in the domain $L$.